OpenStudy (anonymous):

The sum of three consecutive numbers is twelve less than six times the lowest number. What are the three numbers? A. 3, 4, 5 B. 10, 11, 12 C. 2, 3, 4 D. 5, 6, 7

OpenStudy (anonymous):

5,6,7

OpenStudy (anonymous):

5+6+7=18 5*6=30 30-12=18

OpenStudy (texaschic101):

3 consecutive numbers (x) , (x + 1), (x + 2) x + x + 1 + x + 2 = 6x - 12 combine like terms 3x + 3 = 6x - 12 3x - 6x = -12 - 3 -3x = - 15 x = - 15/-3 x = 5 x + 1 = 6 x + 2 = 7 @jesusalejandro94 ...is correct....your numbers are 5,6,7